# Neural Surface Detection for Unsigned Distance Fields

### v1

@@ 4.1 Neural Surface Detection and Marching Cubes
On the MGN dataset we use the training split, consisting of 300 garments, at the four resolutions $n$ 32, 64, 128, or 256. For Shapenet-Cars, due to the greater complexity of the models, we use 64, 128, 256, and 512. We provide qualitative results in Fig. 3, and report quantitative ones in Tab. 1. They are consistent with those observed on ABC. Our method, which is only trained on 80 shapes from ABC, generalizes well to very different geometric structures. The advantage of our method is especially evident at lower resolutions because, at higher resolutions, the surfaces become easier to mesh. Nevertheless, as can be seen in the middle rows of Fig. 3, our method produces cleaner meshes than the baselines on ShapeNet-Cars even at high resolution, and produces more precise borders on MGN, which can be smoothed with a postprocessing step. The difference in reconstruction accuracy is particularly visible in detailed areas of cars, such as tires, rooftops and front hoods. The only exception is DCUDF, which produces clean and smooth meshes but fails to reconstruct parts of the shapes, likely due to the cutting algorithm.
For all methods, however, there remain holes and artifacts in the reconstructed meshes from Shapenet-Cars, mainly due to the many details and the presence of T-junctions, and uneven borders in MGN, with the exception of DCUDC thanks to its integrated smoothing optimization. The problem is less severe at higher resolutions but cannot be handled by Marching Cubes during triangulation, due to its inability to place vertices inside the cells. In such cases, switching from MC to DC is particularly beneficial, as shown in the following section.
@@ A.6 Execution times
A.7 Additional qualitative results

### v2

@@ 4.1 Neural Surface Detection and Marching Cubes
On the MGN dataset we use the training split, consisting of 300 garments, at the four resolutions $n$ 32, 64, 128, or 256. For Shapenet-Cars, due to the greater complexity of the models, we use 64, 128, 256, and 512. We provide qualitative results in Fig. 3, and report quantitative ones in Tab. 1. They are consistent with those observed on ABC. Our method, which is only trained on 80 shapes from ABC, generalizes well to very different geometric structures. The advantage of our method is especially evident at lower resolutions because, at higher resolutions, the surfaces become easier to mesh. Nevertheless, as can be seen in the middle rows of Fig. 3, our method produces cleaner meshes than the baselines on ShapeNet-Cars even at high resolution, and produces more precise borders on MGN, which can be smoothed with a postprocessing step. The difference in reconstruction accuracy is particularly visible in detailed areas of cars, such as tires, rooftops and front hoods. The only exception is DCUDF, which produces clean and smooth meshes but fails to reconstruct parts of the shapes, likely due to the cutting algorithm. See supplementary for details.
For all methods, however, there remain holes and artifacts in the reconstructed meshes from Shapenet-Cars, mainly due to the many details and the presence of T-junctions, and uneven borders in MGN, with the exception of DCUDF thanks to its integrated smoothing optimization. The problem is less severe at higher resolutions but cannot be handled by Marching Cubes during triangulation, due to its inability to place vertices inside the cells. In such cases, switching from MC to DC is particularly beneficial, as shown in the following section.
@@ A.6 Execution times
A.7 Additional experiments with DCUDF
Table A.6: Additional experiments on DCUDF. Median L2 Chamfer Distance ×$\times 10^{-5}$ with 200k sample points (CD) and Image Consistency (IC) are reported at varying grid resolutions. The best results are in bold. "DCUDF-ls" is executed with a lower smoothing parameter (500 instead of 2000). "DCUDF-nc" is executed without the cutting step. *Resolution is doubled for experiments with ShapeNet-Cars due to the higher complexity of the shapes.
ABC [15] MGN [3] Cars [5] MGN autodec. Cars autodec.
32* DCUDF [12] 1200 78.0 603 76.0 559 69.7 503 77.0 318 83.7
DCUDF-ls 1200 77.8 596 76.3 506 79.0 419 78.6 307 83.8
DCUDF-nc 572 82.5 484 77.4 194 81.9 413 78.3 97.1 84.9
Ours + MC [16] 19.0 91.8 8.09 92.0 9.89 87.4 8.37 91.9 13.2 87.3
64* DCUDF [12] 291 86.5 155 85.6 169 81.3 92.4 87.3 55.0 87.3
DCUDF-ls 292 86.6 153 85.6 117 85.9 36.9 92.1 82.2 86.1
DCUDF-nc 138 88.8 119 85.7 29.6 88.1 61.6 87.4 14.7 89.4
Ours + MC [16] 4.46 95.4 2.10 95.5 4.47 91.2 2.64 94.7 10.0 88.5
128* DCUDF [12] 44.6 92.5 26.1 92.3 113 87.7 6.34 95.4 355 78.2
DCUDF-ls 42.0 92.2 26.5 92.3 113 87.6 5.47 95.3 330 77.1
DCUDF-nc 26.1 93.8 20.0 92.3 9.48 91.2 2.29 95.4 69.6 82.3
Ours + MC [16] 2.52 97.4 1.40 97.1 3.13 94.2 2.06 95.2 14.5 87.9
256* DCUDF [12] 11.3 95.0 8.09 94.8 104 89.5 3.40 91.1 1930 36.0
DCUDF-ls 11.5 95.0 8.11 92.3 127 89.5 7.44 90.6 2250 34.0
DCUDF-nc 8.50 96.0 6.14 94.9 4.58 93.3 4.13 91.1 626 43.8
Ours + MC [16] 2.33 98.2 1.33 97.9 3.03 95.8 2.02 94.8 62.2 83.8
We report in Tab A.6 additional experiments with the DCUDF[12] pipeline. In particular, as suggested by the authors, we experiment with a lower smoothing aggressiveness which benefits lower resolutions, denoted as "DCUDF-ls", and with skipping the cutting step of the algorithm, denoted as "DCUDF-nc". The cutting step ensures that only one of the two collapsed surfaces is kept, however it can fail to cut the surface correctly, leading to missing areas of the object as seen in the main paper. Removing such step, however, makes the surface double layered, which may not be suitable for downstream tasks. The increased accuracy compared to the experiment in the main paper comes at the cost of additional parameter tuning. Overall, similar conclusions as in the main paper apply.
A.8 Additional qualitative results